OpenStudy (anonymous):

An equation is shown below: 5(2x - 8) + 15 = -15 Write the steps you will use to solve the equation and explain each step.

OpenStudy (anonymous):

1) distribute: 10x-40+15=-15 2)combine like terms: 10x-10=0 3)solve for x: x=10/10

OpenStudy (bloomlocke367):

What do you need help with @AliSurfCat? is there something specific you don't understand?